# UniForm: A Unified Multi-Task Diffusion Transformer for Audio-Video Generation

### v3

@@ 3.3.2. Multitask Modeling
After obtaining the joint representation of vision as well as audio as input, we adopt STDiT3 (Zheng et al., 2024) blocks to progressively integrate information from both spatial and temporal domains. In order to integrate textual information, cross attention mechanism is applied in both STDiT3 blocks, which can be seen in Figure 2. Noted that due to spatial limitations, we only presented the spatial version of STDiT3. The temporal version of STDiT3, on the other hand, replaced the spatial attention mechanism within the module with a temporal attention mechanism, while keeping all other settings consistent.
3.3.3. Video & Audio Latent Decoding
Once we obtain the final output from the DiT blocks, we utilize multimodal unpatchify to derive the predicted noise and variance for both the audio and video. During training, the predicted noise is used to compute the loss. During inference, the DiT gradually reduces the noise, ultimately generating latent representation with minimal noise in the final diffusion time step. The latent representation is then divided and reshaped into two distinct modal latent forms, corresponding to the audio and video. Using the decoders of pre-trained VAEs, the latent features of both audio and video are simultaneously reconstructed into generated video frames and audio Mel-spectrograms. Subsequently, these Mel-spectrograms are further converted into audio waveforms using the pre-trained HiFi-GAN (Kong et al., 2020).
3.3.4. Video & Audio Generation Loss
@@
(10) $\mathcal{L}_{a}=\left\|\epsilon^{a}_{t}-Mask_{a}(\epsilon_{\theta}(z^{a}_{t},z^{v}_{t},c))\right\|_{2}^{2},$
@@
(11) $\mathcal{L}_{v}=\left\|\epsilon^{v}_{t}-Mask_{v}(\epsilon_{\theta}(z^{a}_{t},z^{v}_{t},c))\right\|_{2}^{2},$
@@
(12) $\mathcal{L}_{av}=\mathcal{L}_{a}+\mathcal{L}_{v},$
@@
Table 1. Comparison of different methods for V2A task.
@@
      VGGSound       SpecVQGAN (Iashin and Rahtu, 2021)       5.42       31.69       5.23       3.37       0.417
      Diff-Foley (Luo et al., 2024)       4.72       23.94       11.11       3.38       0.386
      V-AURA (Viertola et al., 2025)       2.88       14.80       10.08       2.42       0.366
      VATT (Liu et al., 2024a)       2.77       10.63       11.90       1.48       -
      Frieren (Wang et al., 2024d)       1.34       11.45       12.25       2.73       0.422
      FoleyCrafter (Zhang et al., 2024)       2.51       16.24       15.68       2.30       0.403
@@
Figure 3. Compared with FoleyCrafter in V2A generation on the VGGSound dataset. Our method can generate more accurate prosody and richer high-frequency details.
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets
The training datasets used in this work include VGGSound (Chen et al., 2020), Landscape (Lee et al., 2022), AIST++ (Li et al., 2021), AudioSet-balance (Gemmeke et al., 2017) and AudioSet-Strong (Hershey et al., 2021). VGGSound is an extensive single-label audio-visual dataset comprising more than 200,000 videos for 310 audio classes. It is used for various tasks such as audio classification, multi-modal classification, and zero-shot audio classification. Landscape is a high-fidelity dataset that encompasses video and audio streams, highlighting nine varied natural scenes, including but not limited to raining, splashing water, thunder, and underwater bubbling. AIST++ is a dedicated subset constructed from the AIST dance dataset (Tsuchida et al., 2019), containing 1,020 dance motion sequences spanning 10 distinct dance genres, with a total duration of 5.2 hours. To enhance visual presentation, all videos undergo standardized processing through center-cropping techniques, being uniformly resized to a resolution of 1024×1024 pixels. AudioSet-balance consists of 22,176 segments selected from the AudioSet (Gemmeke et al., 2017), with each class having at least 59 samples. AudioSet-Strong is another subset of AudioSet, which involves approximately 67,000 segments with frame-level annotations (with a resolution of 0.1 seconds). We adhered to the settings in (Hayakawa et al., 2025) for the training split.
Table 2. Comparison of different methods for A2V task.
@@ 4.1.1. Datasets
        Landscape         MM-Diffusion (Ruan et al., 2023)         922         2.85         0.410
        TempoToken (Yariv et al., 2024)         784         4.49         0.540
        Sound-guided Video Generation (Lee et al., 2022)         544         1.16         -
        TPoS (Jeong et al., 2023)         421         1.49         -
@@
Figure 4. Generated samples in the A2V task on the Landscape dataset.
4.1.2. Implementation
@@
For data preprocessing, we resample the videos to 17 fps and then resize them to a resolution of $256\times 256$, and we resample the audios at 16 kHz. Then, we truncate the first 4s of the video and audio samples as the input for VAEs. The pre-trained VAEs from Open-Sora (Zheng et al., 2024) and AudioLDM (Liu et al., 2023) are used to encode/decode videos and audios, respectively. We use pLLaVA (Xu et al., 2024), a Visual Language Model (VLM), to generate text descriptions  [^2] Given that audio inherently has a lower information density than video, this limitation constrains the capability of existing audio language models to accurately characterize audio content. Additionally, inconsistent audio-visual descriptions may hinder the model’s ability to effectively learn audiovisual synchronization. Notably, detailed video captions often inherently encompass attributes of sound-producing objects, including target features, motion patterns, and intensity levels. For these reasons, we exclusively adopt video captions as textual conditions without incorporating audio captions.. As for the Landscape and AIST++ dataset, we use its class labels as captions. Unless otherwise stated, the subsequent sections assume text use as the default condition. Our DiT model utilizes pre-trained weights from image generation (Chen et al., 2024). We set the batch size to 32 and conducted 230 epochs of iterations with a constant learning rate of $1\times 10^{-4}$ using the HybirdAdam optimizer. Linear warmup is adopted as the learning rate scheduling strategy, and the warmup step is set to 1000. In the inference stage, the number of inference steps for each task is uniformly set to 30, and the value of classifier-free guidance is set to 5.
4.1.3. Evaluation metrics
@@
For evaluating video generation, we adpot the Frechet Video Distance (FVD), Kernel Video Distance (KVD) and Inception Score (IS). For audio generation, our evaluation relies on the metrics Frechet Audio Distance (FAD), Frechet Distance (FD), kullback–leibler divergence (KL) and IS. Additionally, we utilize the AV-align (Xing et al., 2024) metric to assess the synchronization between generated audio and video.
4.2. Results on Video to Audio Generation
Table 3. Comparison of different methods for T2AV task.
@@ 4.2. Results on Video to Audio Generation
MM-Diffusion (Ruan et al., 2023) 10.61 186 9.21 0.261 10.58 98 18.90 0.273
AV-DiT (Wang et al., 2024c) 11.17 172 15.41 - 10.17 68 21.01 -
MMDisCo (Hayakawa et al., 2025) 5.52 405 - - 2.17 450 - -
UniForm (ours) 2.35 181 9.45 0.308 1.23 76 17.24 0.288
Figure 5. Generated samples in the T2AV task on the Landscape dataset.
Table 1 lists the comparison results of UniForm with some recent approaches on video to audio generation. As can be observed from the table, our approach outperforms most baselines across the majority of metrics. Specifically, it achieved the top rankings in the FAD and FD metrics with scores of 1.3 and 6.21, respectively, surpassing Frieren (Wang et al., 2024d) and VATT (Liu et al., 2024a), which ranked second in their respective categories. The two methods achieved FAD scores of 1.34 and 2.77, and FD scores of 11.45 and 10.63, respectively. Additionally, our model secured the second position in the IS metric with a score of 15.43, closely followed by FoleyCrafter (Zhang et al., 2024), which ranked first with an IS score of 15.68. Despite exhibiting merely average performance in KL, UniForm achieved the highest score in the AV-align metric. This result demonstrates that our approach effectively enhances the alignment between video and audio. It is noteworthy that other baseline methods are only applicable to the task of V2A. This highlights that our multi-task model can generate audio that is highly relevant to the video content, with a quality that rivals the best current V2A approaches.
Figure 3 presents a visual comparison between the mel spectrograms of the audio generated by our method and that produced by FoleyCrafter (Zhang et al., 2024). Compared to FoleyCrafter, our method exhibits higher visual correlation with the ground truth, especially noticeable in the first half of the mel spectrogram where FoleyCrafter lacks certain elements. Additionally, our approach captures more high-frequency details.
Figure 6. Generated samples in the T2AV task on the AIST++ dataset.
Figure 7. Generated two challenging samples on the VGGSound dataset.
4.3. Results on Audio to Video Generation
@@
The comparison results of V2A generation on the Landscape dataset are presented in Table 2. Specifically, our approach achieves the lowest FVD score of 219, demonstrating its ability to generate videos with the highest quality. Meanwhile, in terms of IS, which measures content diversity and quality, our approach also leads, achieving a score of 4.61. This score showcases the strength of our method in generating varied content. Although our method attains an AV-align score of 0.497, slightly lower than the top-ranked TempoToken’s 0.54, our approach overall excels in audio-to-video generation. It proves that UniForm can significantly improve the quality of generated videos while enhancing audio-visual synchronization. Figure 4 presents four representative examples generated by our method on the Landscape dataset.
4.4. Results on Joint Audio-Video Generation
Table 3 lists the comparison results of different methods for joint video and audio generation on the Landscape dataset and AIST++ dataset. As shown in the table, when evaluated on the Landscape dataset, the UniForm algorithm achieves top rankings in both the FAD and AV-align metrics. Additionally, UniForm secures second-place performance in the FVD and KVD metrics, slightly trailing AV-DiT and MM-Diffusion, respectively. For the AIST++ dataset, our method demonstrates leading performance across the FAD, KVD, and AV-align metrics, while exhibiting a slight lag in the FVD metric compared to the AV-DiT approach. Figures 5 and 6 showcase some generation examples of our method on the Landscape dataset and AIST++ dataset, respectively.
@@ 4.4. Results on Joint Audio-Video Generation
Table 4. Comparison with non-unified method across three tasks.
@@
4.5. Comparison with Non-unified Method
Table 4 compares the performance of the UniForm method with non-uniform methods, namely Seeing&Hearing (Xing et al., 2024), across various tasks. Seeing&Hearing employs corresponding distinct pre-trained models when handling different tasks. We evaluate models’ performance on the VGGSound dataset for the V2A and A2V tasks, while for the T2AV task, the Landscape dataset was adopted. This is consistent with the eval datasets used by Seeing&Hearing. It can be observed that our method surpasses the comparative approach across all metrics and tasks, with the exceptions of KL in the V2A task, AV-align in the A2V task, and KVD in the T2AV task. The results highlight the superiority of adopting a unified backbone approach.
4.6. Ablations
4.6.1. The Impact of Text Prompts
Table 5 demonstrates the influence of text on both V2A and A2V tasks. The V2A task was evaluated on the landscape dataset, while the A2V task was conducted on the VGGSound dataset. According to the table, using text prompts can effectively improve all metrics of all tasks, except for the AV-align in the V2A task, where the scores before and after using text prompts are almost identical. Overall, introducing text conditions has noticeably improved the generation performance of both tasks.
Table 5. The influence of text on V2A and A2V tasks.
@@ 4.6.1. The Impact of Text Prompts
A2V 545 42.96 3.01 0.329
219 14.31 4.61 0.497
4.6.2. The Improvement of Alignment in Audio-Visual Joint Generation Compared to Single-modal Generation
Table 6 demonstrates the improvement in alignment on three datasets achieved by the audio-visual joint generative model compared to unimodal generative models, i.e., Unimodal. Unimodal refers to a model where audio and video are independently generated under text guidance, sharing the same architecture as UniForm. As shown in the table, UniForm achieves higher AV-align scores than the independent audio and video generation models across all three datasets. This demonstrates that the proposed model significantly enhances alignment performance between audio and visual modalities, further underscoring the importance of employing a joint generation strategy.
Table 6. The Improvement of Alignment in Audio-Visual Joint Generation Compared to Single-modal Generation.
@@ 4.6.2. The Improvement of Alignment in Audio-Visual Joint Generation Compared to Single-modal Generation
UniForm (Ours) 0.374 0.308 0.288
5. Conclusions
@@
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.

### v5

@@ III-C2 Multitask Modeling
After obtaining the joint representation of vision as well as audio as input, we adopt STDiT3 [8] blocks to progressively integrate information from both spatial and temporal domains. In order to integrate textual information, cross attention mechanism is applied in both STDiT3 blocks, which can be seen in Figure 2. Noted that due to spatial limitations, we only presented the spatial version of STDiT3. The temporal version of STDiT3, on the other hand, replaced the spatial attention mechanism within the module with a temporal attention mechanism, while keeping all other settings consistent.
III-C3 Video & Audio Latent Decoding
Once we obtain the final output from the DiT blocks, we utilize multimodal unpatchify to derive the predicted noise and variance for both the audio and video. During training, the predicted noise is used to compute the loss. During inference, the DiT gradually reduces the noise, ultimately generating latent representation with minimal noise in the final diffusion time step. The latent representation is then divided and reshaped into two distinct modal latent forms, corresponding to the audio and video. Using the decoders of pre-trained VAEs, the latent features of both audio and video are simultaneously reconstructed into generated video frames and audio Mel-spectrograms. Subsequently, these Mel-spectrograms are further converted into audio waveforms using the pre-trained HiFi-GAN [36].
III-C4 Video & Audio Generation Loss
@@
$\mathcal{L}_{a}=\left\|\epsilon^{a}_{t}-Mask_{a}(\epsilon_{\theta}(z^{a}_{t},z^{v}_{t},c))\right\|_{2}^{2},$ (10)
@@
$\mathcal{L}_{v}=\left\|\epsilon^{v}_{t}-Mask_{v}(\epsilon_{\theta}(z^{a}_{t},z^{v}_{t},c))\right\|_{2}^{2},$ (11)
@@
$\mathcal{L}_{av}=\mathcal{L}_{a}+\mathcal{L}_{v},$ (12)
@@
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
The training datasets used in this work include VGGSound [37], Landscape [28], AIST++ [38], AudioSet-balance [39] and AudioSet-Strong [40]. VGGSound is an extensive single-label audio-visual dataset comprising more than 200,000 videos for 310 audio classes. It is used for various tasks such as audio classification, multi-modal classification, and zero-shot audio classification. Landscape is a high-fidelity dataset that encompasses video and audio streams, highlighting nine varied natural scenes, including but not limited to raining, splashing water, thunder, and underwater bubbling. AIST++ is a dedicated subset constructed from the AIST dance dataset [41], containing 1,020 dance motion sequences spanning 10 distinct dance genres, with a total duration of 5.2 hours. To enhance visual presentation, all videos undergo standardized processing through center-cropping techniques, being uniformly resized to a resolution of 1024×1024 pixels. AudioSet-balance consists of 22,176 segments selected from the AudioSet [39], with each class having at least 59 samples. AudioSet-Strong is another subset of AudioSet, which involves approximately 67,000 segments with frame-level annotations (with a resolution of 0.1 seconds). We adhered to the settings in [15] for the training split.
TABLE I: Comparison of different methods for V2A task.
@@ IV-A1 Datasets
      VGGSound       SpecVQGAN [21]       5.42       31.69       5.23       3.37       0.417
      Diff-Foley [23]       4.72       23.94       11.11       3.38       0.386
      V-AURA [25]       2.88       14.80       10.08       2.42       0.366
      VATT [26]       2.77       10.63       11.90       1.48       -
      Frieren [27]       1.34       11.45       12.25       2.73       0.422
      FoleyCrafter [24]       2.51       16.24       15.68       2.30       0.403
@@
TABLE II: Comparison of different methods for A2V task.
@@
        Landscape         MM-Diffusion [12]         922         2.85         0.410
        TempoToken [30]         784         4.49         0.540
        Sound-guided Video Generation [28]         544         1.16         -
        TPoS [29]         421         1.49         -
@@
IV-A2 Implementation
@@
For data preprocessing, we resample the videos to 17 fps and then resize them to a resolution of $256\times 256$, and we resample the audios at 16 kHz. Then, we truncate the first 4s of the video and audio samples as the input for VAEs. The pre-trained VAEs from Open-Sora [8] and AudioLDM [20] are used to encode/decode videos and audios, respectively. We use pLLaVA [42], a Visual Language Model (VLM), to generate text descriptions  [^2] Given that audio inherently has a lower information density than video, this limitation constrains the capability of existing audio language models to accurately characterize audio content. Additionally, inconsistent audio-visual descriptions may hinder the model’s ability to effectively learn audiovisual synchronization. Notably, detailed video captions often inherently encompass attributes of sound-producing objects, including target features, motion patterns, and intensity levels. For these reasons, we exclusively adopt video captions as textual conditions without incorporating audio captions.. As for the Landscape and AIST++ dataset, we use its class labels as captions. Unless otherwise stated, the subsequent sections assume text use as the default condition. Our DiT model utilizes pre-trained weights from image generation [43]. We set the batch size to 32 and conducted 230 epochs of iterations with a constant learning rate of $1\times 10^{-4}$ using the HybirdAdam optimizer. Linear warmup is adopted as the learning rate scheduling strategy, and the warmup step is set to 1000. In the inference stage, the number of inference steps for each task is uniformly set to 30, and the value of classifier-free guidance is set to 5.
Figure 3: Compared with FoleyCrafter in V2A generation on the VGGSound dataset. Our method can generate more accurate prosody and richer high-frequency details.
Figure 4: Generated samples in the A2V task on the Landscape dataset.
IV-A3 Evaluation metrics
@@
For evaluating video generation, we adpot the Frechet Video Distance (FVD), Kernel Video Distance (KVD) and Inception Score (IS). For audio generation, our evaluation relies on the metrics Frechet Audio Distance (FAD), Frechet Distance (FD), kullback–leibler divergence (KL) and IS. Additionally, we utilize the AV-align [17] metric to assess the synchronization between generated audio and video.
IV-B Results on Video to Audio Generation
Table I lists the comparison results of UniForm with some recent single-task methods on V2A generation, with the evaluation dataset being the VGGSound test set. As can be observed from the table, our approach outperforms most baselines across the majority of metrics. Specifically, it achieved the top rankings in the FAD and FD metrics with scores of 1.3 and 6.21, respectively, surpassing Frieren [27] and VATT [26], which ranked second in their respective categories. The two methods achieved FAD scores of 1.34 and 2.77, and FD scores of 11.45 and 10.63, respectively. Additionally, our model secured the second position in the IS metric with a score of 15.43, closely followed by FoleyCrafter [24], which ranked first with an IS score of 15.68. Despite exhibiting merely average performance in KL, UniForm achieved the highest score in the AV-align metric. This result demonstrates that our approach effectively enhances the alignment between video and audio. It is noteworthy that other baseline methods are only applicable to the task of V2A. This highlights that our multi-task model can generate audio that is highly relevant to the video content, with a quality that rivals the best current V2A approaches.
Figure 3 presents a visual comparison between the mel spectrograms of the audio generated by our method and that produced by FoleyCrafter [24]. Compared to FoleyCrafter, our method exhibits higher visual correlation with the ground truth, especially noticeable in the first half of the mel spectrogram where FoleyCrafter lacks certain elements. Additionally, our approach captures more high-frequency details.
TABLE III: Comparison of different methods for T2AV task.
@@ IV-B Results on Video to Audio Generation
MM-Diffusion [12] 10.61 186 9.21 0.261 10.58 98 18.90 0.273
MM-LDM [13] 9.1 77 3.20 - 10.2 55 8.20 -
AV-DiT [16] 11.17 172 15.41 - 10.17 68 21.01 -
MMDisCo [15] 5.52 405 - - 2.17 450 - -
UniForm (ours) 2.41 164 9.05 0.305 1.27 62 17.24 0.283
Figure 5: Generated samples in the T2AV task on the Landscape dataset.
Figure 6: Generated samples in the T2AV task on the AIST++ dataset.
Figure 7: Generated two challenging samples on the VGGSound dataset.
IV-C Results on Audio to Video Generation
@@
Table II presents the comparison between UniForm and A2V-specific baselines on the Landscape dataset. Specifically, our approach achieves the lowest FVD score of 219, demonstrating its ability to generate videos with the highest quality. Meanwhile, in terms of IS, which measures content diversity and quality, our approach also leads, achieving a score of 4.61. This score showcases the strength of our method in generating varied content. Although our method attains an AV-align score of 0.497, slightly lower than the top-ranked TempoToken’s 0.54, our approach overall excels in audio-to-video generation. It proves that UniForm can significantly improve the quality of generated videos while enhancing audio-visual synchronization. Figure 4 presents four representative examples generated by our method on the Landscape dataset.
IV-D Results on Joint Audio-Video Generation
Table III lists the comparison results of different methods for joint video and audio generation on the Landscape dataset and AIST++ dataset. In evaluations on the Landscape and AIST++ datasets, the UniForm algorithm outperforms MM-Diffusion, AV-DiT, and MMDisCo across various metrics for both audio and video generation. Compared to MM-LDM, our method performs worse on video generation metrics across the two datasets but shows significant superiority in audio generation. The gap in video generation may be due to MM-LDM being trained separately on the two datasets with dataset-specific training (unlike our approach based on large-scale data training), resulting in generated distributions that are more closely aligned with the original data distributions. Additionally, our method also outperforms the comparative methods in audio-video consistency. Figures 5 and 6 showcase some generation examples of our method on the Landscape dataset and AIST++ dataset, respectively.
@@ IV-D Results on Joint Audio-Video Generation
TABLE IV: Comparison with non-unified method across three tasks.
@@
IV-E Comparison with Non-unified Method
Table IV compares the performance of the UniForm method with non-uniform methods, namely Seeing&Hearing [17], across various tasks. Seeing&Hearing employs corresponding distinct pre-trained models when handling different tasks. We evaluate models’ performance on the VGGSound dataset for the V2A and A2V tasks, while for the T2AV task, the Landscape dataset was adopted. This is consistent with the eval datasets used by Seeing&Hearing. It can be observed that our method surpasses the comparative approach across all metrics and tasks, with the exceptions of KL in the V2A task, AV-align in the A2V task, and KVD in the T2AV task. The results highlight the superiority of adopting a unified backbone approach.
TABLE V: The influence of text on V2A and A2V tasks.
@@ IV-E Comparison with Non-unified Method
       A2V        ✗        545        42.96        3.01        0.329
       ✓        219        14.31        4.61        0.497
TABLE VI: The Improvement of Alignment in Audio-Visual Joint Generation Compared to Single-modal Generation.
@@
UniForm (Ours) 0.374 0.305 0.283
IV-F Ablations
IV-F1 The Impact of Text Prompts
Table V demonstrates the influence of text on both V2A and A2V tasks. The V2A task was evaluated on the landscape dataset, while the A2V task was conducted on the VGGSound dataset. According to the table, using text prompts can effectively improve all metrics of all tasks, except for the AV-align in the V2A task, where the scores before and after using text prompts are almost identical. Overall, introducing text conditions has noticeably improved the generation performance of both tasks.
IV-F2 The Improvement of Alignment in Audio-Visual Joint Generation Compared to Single-modal Generation
Table VI demonstrates the improvement in alignment on three datasets achieved by the audio-visual joint generative model compared to unimodal generative models, i.e., Unimodal. Unimodal refers to a model where audio and video are independently generated under text guidance, sharing the same architecture as UniForm. As shown in the table, UniForm achieves higher AV-align scores than the independent audio and video generation models across all three datasets. This demonstrates that the proposed model significantly enhances alignment performance between audio and visual modalities, further underscoring the importance of employing a joint generation strategy.
V Conclusions